This is the fourth time i ask the same question and nobody can give me the right answer, so if you dont know how to solve it don't answer it becasue i cant waste any more of my question on this one question.

This is an integrated rate law equation problem.

This is the fourth time i ask the same ques

The last tutor used an equation with 2.303 i do not know where they come up with this number, so if you use it in solving this problem please explain where it comes from.

first order rate law:

In[A]=-kt+In[A initial]
Answer #1

A)

k = 87 s-1

initial concentration = Ao = 2.0 M

for first order reaction:

k = 1/t * ln (Ao /At)

87 = 1/ 0.010 * ln (2 / At)

0.87 = ln (2 /At)

At = 0.84 M

concentration after 0.010 s = 0.84 M

B)

initial = 2.0 M

remaining = 0.84 M

decomposed = 2-0.84 = 1.16 M

fraction of decompsoed concetration = 1.16 / 2 .0

                                                       = 0.58